OpenStudy (anonymous):

In 1981 108.5 million kg of butter were producet in England. In 1990 93.3 million kg of butter were produced in England. How many percent of the production of butter has decreased in England from 1981 to 1990? If we assume that the decrease has been even over whole periods, how big has the annual decrease in percentage been? How big a percentage of the production of butter in 1981 make up of the production of butter in 1990? Make a mathematical model that describes the production of butter for 1981 to 1990?

OpenStudy (cruffo):

How many percent of the production of butter has decreased in England from 1981 to 1990? This seems to be asking for the percent decrease % decrease = (old value - new value)/old value = 14% Not sure what questions 2 and 3 are asking. Make a mathematical model that describes the production of butter for 1981 to 1990? Since you only have two points of data, assume you can use a line: \(m=\frac{y_2-y_1}{x_2-x_1}\) and then \(y - y_1 = m(x-x_1)\) Let \(x_2 = 1990\) and \(y_2 = 93.3\) and \(x_1 = 1981\) and \(y_2 = 108.5\)

OpenStudy (anonymous):

Question 2, I will reformulate, it's hard because English is not my first language :) We assume that the decrease have been regular, how big has the annual decrease in percent been?

OpenStudy (cruffo):

I'm sorry, I'm still not sure what is being asked.. I get the first part. "If production of butter has been decreasing by 14% for several years now, " Just not sure what the second part means: "how big" It may be that they just want you to state the percent as a proportion, as in "they are producing 14 million kg less every year for every 100 million kg produced" or something like that.